OpenStudy (anonymous):

Determine the sum of the given probabilities. Are the events complementary? P(A) = ; P(B) = A. The events are complementary. B. The events are not complementary. C. The events are complementary. D. The events are not complementary.

OpenStudy (anonymous):

p(a)=3/7 p(b)=2/7

OpenStudy (anonymous):

A.5/14 the events are complementary. B.5/14 The events are not complementary. C.5/7 The events are complementary. D.5/7 The events are not complementary.

OpenStudy (anonymous):

@dan815 @jigglypuff314

OpenStudy (anonymous):

@poopsiedoodle @EclipsedStar @WingedArtist @rose18920 @kiamousekia @oblivion20 @AwesomeSays3! @andrewhaze @Ria23 @EmmanuelRocksYou

OpenStudy (anonymous):

i dont know this :/

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

In the misty morning fog with Our, our hearts a-thumping And you, my brown-eyed girl, You, my brown-eyed girl.

OpenStudy (anonymous):

ummmmmmmmmm ok @EmmanuelRocksYou

OpenStudy (anonymous):

@Chiko_1278